# Class Distance Weighted Cross Entropy Loss for Classification of Disease Severity

### v1

@@ Methods
Results
We performed experiments for Ulcerative Colitis classification using the dataset detailed in the previous section. The performance results are complemented by an explainability analysis utilizing t-SNE plots. Additionally, we gathered feedback from domain experts to evaluate whether our loss function contributes to the model’s Class Activation Map (CAM) in better aligning with professional opinion.
Assessment of Ulcerative Colitis Classification Performance
We have employed three deep learning models and six different loss functions, reporting results against various loss functions. The performance results are reported in Table 1 and 2 for multi-class (MES) classification and binary score (remission) classification tasks, respectively.
@@ Results
Loss Function ResNet18 Inception-v3 MobileNet-v3-Large
QWK Cross-Entropy 0.8296 ± 0.014 0.8360 ± 0.011 0.8302 ± 0.015
@@
Classifying disease severity into remission and non-remission holds clinical significance. Consequently, we mapped the results in Table 2 into remission classification without training a new model. CDW-CE loss function demonstrates the best performance by achieving the highest Kappa, F1-Score, and accuracy metrics. Contrary to multi-class classification, the CE loss function often performs worst among other loss functions while achieving better scores than CORN and CO2 methods for binary classification. Similar to multi-class classification, MSE exhibits better performance than other loss functions (CORN, CO2, HO2, and CE).
@@
Loss Function ResNet18 Inception-v3 MobileNet-v3-Large
Kappa Cross-Entropy 0.8077 ± 0.023 0.8074 ± 0.021 0.8122 ± 0.018
@@
(b) Mean confusion matrix for remission classification.
@@
Discussion
Explainability Analysis using t-SNE plots
In this section, we use t-SNE plots [24] to visually assess the feature extraction capability of our method by examining high-dimensional data. t-SNE effectively reveals clusters and latent patterns in the data. To complement this, we analyze silhouette scores, which quantify clustering quality by measuring intra-class cohesion and inter-class separation. A higher silhouette score indicates better clustering. While t-SNE plots are helpful for visual insights, silhouette scores provide an objective comparison of different approaches.
(a) ResNet18 with CE.
Silhouette score: 0.121
(b) ResNet18 with CDW-CE.
Silhouette score: 0.222
Figure 5: t-SNE embeddings of features obtained from ResNet18 model trained with a) CE b) CDW-CE
Figures 5(a) and 5(b) display the t-SNE plots illustrating the embeddings of our models trained on LIMUC using CE and CDW-CE, respectively. The plots reveal a notable separation between MES 0 and MES 3 classes compared to MES 1 and MES 2 classes. However, these plots do not distinctly indicate the relative embedding qualities. Therefore, we employed the Silhouette Score to quantitatively compare the feature extraction ability of the two loss functions. For $N$ samples, the average Silhouette Score is calculated as shown in Eq. 5.
$S=\frac{1}{N}\sum_{i=1}^{N}\frac{b_{i}-a_{i}}{\max(a_{i},b_{i})}$ (5)
where $a_{i}$ is the mean distance from the $i^{th}$ sample to other samples in the same cluster, and $b_{i}$ is the mean distance from the $i^{th}$ sample to samples in the nearest cluster.
The Silhouette scores for models trained with CE and CDW-CE were calculated as 0.121 and 0.222, respectively. This suggests that models trained with CDW-CE exhibit a better feature extraction ability compared to those trained with the standard CE loss function.
Explainability Analysis: Class Activation Maps
@@ Discussion
The addition of the margin term improves the CDW-CE loss function in terms of classification performance as demonstrated by a ResNet18 architecture trained with and without margin term CDW-CE loss functions with different $\alpha$ values in Figure 8. Here, we have approached the margin term as a new hyperparameter and tuned it according to the QWK performance. With this tuning, we have achieved improved results for different margin values corresponding to different power terms ($\alpha$). The $\alpha=1$ is the basic CDW-CE loss function with no class distance weights, and the inclusion of the margin term improves the results in this baseline form. It is important to note that the extra hyperparameter tuning effort is required to obtain a possibly better result than default values for the margin value for different tasks and datasets.
@@
Adding a margin term effectively improves all three models used in this work in terms of QWK score.
@@
Loss function ResNet18 Inception-v3 MobileNet-v3-large
CDW-CE 0.8568 0.8678 0.8588
CDW-CE w/ margin 0.8600 ($m$=0.05) 0.8719 ($m$=0.025) 0.8617 ($m$=0.0025)
Conclusions
In this work, we have proposed a new loss function, Class Distance Weighted Cross-Entropy (CDW-CE), designed specifically for ordinal classification tasks. Our experiments on the LIMUC dataset demonstrated that ordinal loss functions perform better than the widely used non-ordinal loss function CE and other ordinal loss functions. Moreover, the proposed CDW-CE method outperformed the other methods across various performance metrics and architectures. Additionally, an analysis of remission scores indicated that CDW-CE surpasses other ordinal loss functions (CO2, HO2, CORN) in remission scores.
To demonstrate the encoding capability of models trained with our approach, we generated t-SNE plots. We calculated corresponding silhouette scores for CDW-CE and CE loss functions, representing ordinal and categorical scenarios. The Silhouette Score comparison revealed that CDW-CE, with higher scores, exhibits superior intra-class grouping and better inter-class separation. Furthermore, feedback from domain experts based on CAM visualizations indicated that our model outperformed the baseline CE loss function, suggesting improved explainability.
Acknowledgements
@@
Data availability
Labeled Images for Ulcerative Colitis (LIMUC) [7] dataset is a publicly available dataset.
Conflicts of interest

### v2

@@ 3.4 Dataset
4 Results
We performed experiments for Ulcerative Colitis classification using the dataset detailed in the previous section. The performance results are complemented by an explainability analysis utilizing t-SNE and UMAP plots. Additionally, we gathered feedback from domain experts to evaluate whether our loss function contributes to the model’s Class Activation Map (CAM) in better aligning with professional opinion.
4.1 Assessment of Ulcerative Colitis Classification Performance
In this study, we have reported Quadratic Weighted Kappa (QWK), F1 Score, accuracy and Mean Absolute Error (MAE) as well as the receiver operating characteristics (ROC) curves with area under the curve (AUC) for true positive and false positive rates. Since precision and recall metrics are used to calculate the F1 score, they were deemed redundant and were not reported separately.
We have employed three deep learning models and six different loss functions, reporting results against various loss functions. The performance results are reported in Table 2 and 3 for multi-class (MES) classification and binary score (remission) classification tasks, respectively.
@@ 4.1 Assessment of Ulcerative Colitis Classification Performance
Loss ResNet18 Inception-v3 MobileNet-v3-L
QWK CE 0.8296 ± 0.014 0.8360 ± 0.011 0.8302 ± 0.015
@@
Classifying disease severity into remission and non-remission holds clinical significance. Consequently, we mapped the results in Table 3 into remission classification without training a new model. CDW-CE loss function demonstrates the best performance by achieving the highest Kappa, F1-Score, and accuracy metrics. Contrary to multi-class classification, the CE loss function often performs worst among other loss functions while achieving better scores than CORN and CO2 methods for binary classification. Similar to multi-class classification, MSE exhibits second best performance.
@@
Loss ResNet18 Inception-v3 MobileNet-v3-L
Kappa CE 0.8077 ± 0.023 0.8074 ± 0.021 0.8122 ± 0.018
@@
(b) Mean confusion matrix for remission.
@@
Figure 5: ROC curves obtained with the same deep learning architecture trained with different loss functions.
Receiver operating characteristic (ROC) curves measure the trade-off between false positive and true positive rates, with the area under the curve (AUC) indicating a model’s ability to discriminate between classes. Figure 5 presents the ROC curves and corresponding AUC scores for each MES class,based on the ResNet18 architecture trained with various loss functions.
The ROC curves for various loss functions are presented in Figure 5. The results show that CO2 and HO2 exhibit significantly worse performance for the MES 1 class, while achieving similar or better performance for other classes compared to the CE loss function. The model trained with CDW-CE outperforms all others, including ordinal loss functions, by achieving higher AUC scores across all classes. Analyzing the ROC curves for different loss functions reveals that CDW-CE not only improves classification performance in terms of QWK and F1 scores but also improves the balance between false positive and true positive rates.
5 Discussion
5.1 Explainability Analysis using t-SNE plots
In this section, we use t-SNE (Van der Maaten and Hinton, 2008) and UMAP (McInnes et al., 2018) plots to visually assess the feature extraction capability of our method by examining high-dimensional data. UMAP and t-SNE effectively reveals clusters and latent patterns in the data. To complement this, we analyze silhouette scores, which quantify clustering quality by measuring intra-class cohesion and inter-class separation. A higher silhouette score indicates better clustering. While t-SNE and UMAP plots are helpful for visual insights, silhouette scores provide an objective comparison of different approaches.
(a) CE t-SNE mean Silhouette Score: 0.156
(b) CO2 t-SNE mean Silhouette Score: 0.138
(c) HO2 t-SNE mean Silhouette Score: 0.092
(d) CDW-CE t-SNE mean Silhouette Score: 0.176
Figure 6: t-SNE plots obtained with the same deep learning architecture trained with different loss functions.
Figures 6 and 7 display the t-SNE and UMAP plots illustrating the embeddings of our models trained on LIMUC using CE and CDW-CE, respectively. The plots reveal a notable separation between MES 0 and MES 3 classes compared to MES 1 and MES 2 classes.
(a) CE UMAP Silhouette Score: 0.167
(b) CO2 UMAP Silhouette Score: 0.163
(c) HO2 UMAP Silhouette Score: 0.089
(d) CDW-CE UMAP Silhouette Score: 0.201
Figure 7: UMAP plots obtained with the same deep learning architecture trained with different loss functions.
The mean silhouette scores for models trained with various loss functions are presented in Table 4. The results indicate that models trained using CDW-CE achieve higher silhouette scores on both t-SNE and UMAP plots, suggesting better feature extraction capabilities than models trained with other loss functions.
Loss Function t-SNE UMAP
CE 0.156 0.167
CO2 0.138 0.163
HO2 0.092 0.089
CDW-CE 0.176 0.201
Table 4: Mean silhouette scores of t-SNE and UMAP plots for models trained with different loss functions.
5.2 Explainability Analysis: Class Activation Maps
@@ 5.3 Analysis of the Hyperparameters
We analyzed the impact of $\alpha$ by varying it across integer values within the range of 1 to 10. For each model (ResNet18, Inception-v3, and MobileNet-v3-large), we found the best-performing power term values as 5, 6, and 7, respectively. These experiments are summarized in the Figure 10.
The addition of the margin term improves the CDW-CE loss function in terms of classification performance as demonstrated by a ResNet18 architecture trained with and without margin term CDW-CE loss functions with different $\alpha$ values in Figure 11. Here, we have approached the margin term as a new hyperparameter and tuned it according to the QWK performance. With this tuning, we have achieved improved results for different margin values corresponding to different power terms ($\alpha$). The $\alpha=1$ is the basic CDW-CE loss function with no class distance weights, and the inclusion of the margin term improves the results in this baseline form.
@@
Adding a margin term effectively improves all models used in this work in terms of QWK score as shown in Table 5.
@@
CDW-CE ResNet18 Inception-v3 MobileNet-v3-L
no margin 0.8568 0.8678 0.8588
with margin 0.8600 ($m$=0.05) 0.8719 ($m$=0.025) 0.8617 ($m$=0.0025)
6 Conclusions
In this work, we have proposed a new loss function, Class Distance Weighted Cross-Entropy (CDW-CE), designed specifically for ordinal classification tasks. Our experiments on the LIMUC dataset demonstrated that ordinal loss functions perform better than the widely used non-ordinal loss function CE and other ordinal loss functions. Moreover, the proposed CDW-CE method outperformed the other methods across various performance metrics and architectures. With ROC curves and corresponding AUC scores, CDW-CE outperformed other loss functions. Additionally, an analysis of remission scores indicated that CDW-CE surpasses other ordinal loss functions.
To demonstrate the encoding capability of models trained with our approach, we generated t-SNE and UMAP plots. We calculated corresponding silhouette scores for CE, HO2, CO2 and CDW-CE loss functions, representing ordinal and categorical scenarios. The Silhouette Score comparison revealed that CDW-CE, with higher scores, exhibits superior intra-class grouping and better inter-class separation. Furthermore, feedback from domain experts based on CAM visualizations indicated that our model outperformed the baseline CE loss function, suggesting improved explainability.
7 Acknowledgements
@@
8 Author contributions
All authors took part in the conceptualization. Gorkem Polat performed data curation, formal analysis, investigation, validation, visualization and contributed to the methodology, software and revision of the manuscript. Ümit Mert Çağlar contributed to the data curation, software, visualization, preparation of the original draft and editing the manuscript. Alptekin Temizel provided resources, funding acquisition, supervision, project administration and contributed to the methodology, validation and review of the manuscript.
9 Data availability
Labeled Images for Ulcerative Colitis (LIMUC) dataset is a publicly available dataset, accessible at https://zenodo.org/records/5827695.
10 Conflicts of interest